### v1

@@ 2.3.3 Single-snapshot security
A PD scheme $\Pi$ is single-snapshot (SS)-secure iff for any PPT adversary $\mathcal{A}$ which chooses $\ell\leq\mathsf{max}$, passwords $P_{1},\ldots,P_{\ell-1}$, and access patterns $O_{0}$ and $O_{1}$ subject to the constraints outlined in Section 2.3.2, it holds:
@@
  • $D_{0}$ is the disk snapshot resulting from the application of $O_{0}$ to $\Sigma_{0}$, where $\Sigma_{0}\leftarrow\Pi.\mathtt{setup}(P_{1},\ldots,P_{\ell})$ and $P_{\ell}\raisebox{-1.0pt}{$\,\xleftarrow{\raisebox{-1.0pt}{$\scriptscriptstyle\$$}}\,$}\{0,1\}^{\lambda}$.
  • $D_{1}$ is the disk snapshot resulting from the application of $O_{1}$ to $\Sigma_{1}$, where $\Sigma_{1}\leftarrow\Pi.\mathtt{setup}(P_{1},\ldots,P_{\ell-1})$.
@@ 4 The Shufflecake Scheme
In this section we present our Shufflecake scheme, we explain its way of operation, and we provide a security analysis.
@@ 4.1 Design
    Initialize a device: this is done only once, when a new device is first prepared for use in Shufflecake. It consists in overwriting the device with random data, asking the user to provide the number $\ell$ of desired volumes and related passwords, and creating an encrypted header with metadata using this information. In our implementation, this functionality is provided by the init command.
@@
    Instantiate a device: this is the preliminary stage of preparing a Shufflecake-initialized device for use. It consists of reading the user’s provided password, trying to decrypt the device’s header metadata with the derived key, and, if successful, recover information on the available volumes provided. In our implementation this functionality is invisible to the user (it gets executed together with open), so we don’t provide an associated command.
@@
    Open a volume: using the correctly derived volume key, volume-specific metadata is read from the relevant header section. This metadata is used to create a logical device which is presented to the user, and the user’s OS can issue SflcRead and SflcWrite requests to this logical volume. In our implementation, this functionality is provided by the open command.
@@
    Close a previously instantiated device: ephemeral state changes, if present, are written (encrypted) to disk, and then all the open volumes provided by that device are removed from the user’s view. In our implementation, this functionality is provided by the close command.
At its core, Shufflecake is a block indirection layer on top of an encryption layer. Our indirection layer realizes a mechanism which is already a strong improvement over TrueCrypt, since it fixes two of its crucial limitations: it allows multiple volumes, and it is filesystem-independent. Data decryption keys for every volume are derived by a password and other header-derived randomness. Furthermore, the decrypted payload of the header of volume $\ell>1$ also contains a copy of the header decryption key for volume $\ell-1$. This allows to recursively open all volumes present in a device by using a single password, which in turn improves security and user experience, as we will see.
@@
The device’s physical storage space is statically divided into a header section and a data section. The header section is found at the beginning of the disk, and it is composed of a fixed-size device master block (DMB), and $\mathsf{max}$ equal-sized volume headers (each of them comprised of a volume master block (VMB) and a position map), irrespective of how many volumes there are effectively. This mild waste of space is necessary in order to to prevent the adversary from trivially deducing the number of volumes by the size of the device header (which might be possible by analysing the data allocation pattern even when not all volumes are opened). Let us analyse all these sections, starting first with the data.
@@
We distinguish between logical slices, of size $S_{L}$ blocks, and physical slices, of size $S_{P}=S_{L}+\Delta_{S}$, where $\Delta_{S}$ blocks are used to store the encryption IVs for that slice. Logical slices store the data of their respective volumes, while physical slices are used to reserve space on the disk to allocate the encrypted logical slice. A physical slice can either be unallocated (i.e., unclaimed by any volume), or mapped to a single logical slice belonging to some volume $V_{i}$. In the latter case, the mapping from the logical slice index (LSI) $\sigma$ of volume $V_{i}$ to the device’s physical slice index (PSI) $\varPsi$ is given by a function $\mathtt{GetSliceMap}:(i,\sigma)\mapsto\varPsi$ which is computed by simply looking up a position map (basically, a per-volume array). I/O operations to a logical block address $B$ of volume $V_{i}$ are performed through the two interfaces $\mathtt{SflcRead}$ and $\mathtt{SflcWrite}$. We will describe later these two interfaces, as well as the structure of the position map.
@@
The DMB encapsulates all password-related data. It begins with one single KDF salt, shared for all volumes: this salt, combined with a volume $V_{i}$’s password through a KDF, yields the volume’s key-encryption-key ($\mathsf{KEK}_{i}$). Notice that we derive every $\mathsf{KEK}_{i}$ by using just a single global salt for all $\mathsf{max}$ volumes, otherwise we would incur in up to $\mathsf{max}$ expensive different key derivations every time we instantiate a device [^5] This limitation can actually be avoided by using some key derivation tricks, like ‘re-salting’, i.e. using the output of the KDF in combination with a (fast) hash using a per-header salt. We leave this approach for future consideration.. This does not hamper security, as password hashes are never stored on disk, but only used to generate the KEK which is in turn used as a decryption key.
Then come $\mathsf{max}$ DMB cells, each being an authenticated ciphertext (together with the corresponding IV), encrypted with the respective volume’s $\mathsf{KEK}$: the plaintext is itself another cryptographic key, encrypting the volume’s VMB (the volume master key $\mathsf{VMK}_{i}$). This key decoupling allows us to possibly change a volume’s password without having to re-encrypt all its content with a different key. For granularity and consistency, the overall size of the DMB is fixed to be exactly one block. The rest of the DMB, and the unused DMB cells, contain random noise.
@@
The $i$-th volume header is composed of a Volume Master Block (VMB), followed by the encrypted (but not authenticated) position map of volume $V_{i}$. We discuss the position map in the next paragraph. The VMB is a single block containing a (non-authenticated) ciphertext, encrypted with the volume master block’s key $\mathsf{VMK}_{i}$, and the associated IV. The underlying plaintext is composed of the following fields:
  • The volume’s volume-encryption key ($\mathsf{VEK}_{i}$), used for encrypting the actual data section and the position map.
  • The previous volume’s VMB key $\mathsf{VMK}_{i-1}$ (or a random value if $i=1$).
  • The number of slices $\mathtt{NumSlices}$ contained in the device.
@@
The device-specific value $\mathtt{NumSlices}$ defines and fixes the size of the position maps, even in the case that the device is resized; it is replicated across all volume headers in order to be decryptable with any provided password. The presence of $\mathsf{VMK}_{i-1}$ is what allows us to impose a hierarchy on the otherwise-independent volumes (they are all treated equally in the data section). This way, once we open volume $V_{i}$, we can iteratively walk the backwards-linked list induced by this field, and also open volumes $V_{i-1}$ through $V_{1}$. While this approach compromises deniability for a volume $V_{i}$ “in the middle”, once the password to $V_{j}\,,\,j>i$ has been provided, it does not harm deniability as defined by the security game: the volume we want to hide is the last one, $V_{\ell}$, not the middle ones. The usefulness of this approach is discussed in Section 4.2.
@@
Slice maps are arrays $\mathtt{SliceMap}_{i}$ of $\mathtt{NumSlices}$ elements, where every element is a PSI: the index of each element is the LSI mapping to that PSI. Each volume’s slice map is decrypted and loaded in memory when the volume is instantiated; these maps are kept entirely in RAM while the volumes are “live”, and persisted on-disk (encrypted, together with their fresh IVs) when the volumes are closed. Slice maps are stored after the VMBs as equal-size ciphertexts, large enough to address all the possible physical slices. The RAM and disk space requirements are modest: if the underlying device is $N$ blocks large, the size of a slice map is just $O\left(\frac{N}{S_{P}}\log\frac{N}{S_{P}}\right)$, because there are at most $\frac{N}{S_{P}}$ physical slices, each requiring $O\left(\log\frac{N}{S_{P}}\right)$ bits to be indexed. This is in turn due to the choice of addressing the storage space at the slice granularity, instead of the block granularity, which would have entailed a position map of $O(N\log N)$ bits.
@@
We now describe more in detail how Shufflecake operates and explain the rationale behind certain design choices. We start by explaining how we use encryption to secure data on-disk. Then we look at the indirection layer between physical data allocation on disk and data on the logical volumes, i.e., the mechanism which creates the correspondence between physical and logical slices. Finally, we explain how this correspondence is updated when new logical slices are needed as data is written on the logical volumes.
@@
In our case, however, we stick to explicitly random IVs because we want to keep the future possibility to extend Shufflecake with some degree of multi-snapshot security, requiring us to re-encrypt some blocks with a different IV while leaving their content unchanged. For this reason, we use CTR mode instead, and the IV of a block is refreshed at each write for that block, to avoid IV-reuse attacks. This means that all these IVs are stored on-disk.
This strategy introduces a potential performance issue: a naive implementation would translate each logical SflcWrite to a physical bWrite of the corresponding physical data block, plus an additional bRead-update-bWrite of the whole corresponding IV block. This would be very wasteful in terms of I/O overhead, because we only need to update one IV (e.g., 16 bytes for AES-CTR), but we are forced to load and store whole blocks (typically 4096 bytes).
We avoid this problem by caching IV blocks in RAM, in an LRU cache of predefined depth (e.g., 1024 entries). For the performance reasons just discussed, this cache is not write-through. This way, we coalesce possibly many updates of the same IV block (triggered by many logical SflcWrites to the same data block, or by logical SflcWrites to many blocks within the same slice) into just one physical bWrite, thereby lowering the I/O overhead.
For each physical slice, we pack the IVs for the blocks contained therein into $\Delta_{S}$ physical blocks at the beginning of the slice. There is a simple static correspondence between a physical block and the on-disk location of its IV: the $(m-\Delta_{S})$-th block within a physical slice is encrypted by the $m$-th IV within the initial $\Delta_{S}$ blocks. Hence, we assume the existence of functions $\mathtt{LoadIV}$ and $\mathtt{SampleAndStoreIV}$ (with self-explanatory behaviour) which take as input a physical address $B_{\mathsf{phys}}$ and return the corresponding IV. Analogously, we consider the functions $\mathtt{Encrypt}(\textsf{ptxt},\textsf{key};\textsf{IV})$ and $\mathtt{Decrypt}(\textsf{ctxt},\textsf{key};\textsf{IV})$ as acting on blocks according to the implied mode of operation.
@@
Consider a read or write operation to a logical block address $B$ for volume $V_{i}$. There are three possible cases:
@@
    The requested operation (read or write) for volume $i$ happens on a block whose logical address $B$ was previously allocated. In this case we need to map efficiently $B$ to the corresponding physical address $B_{\mathsf{phys}}$.
@@
    It is a $\mathtt{SflcRead}$ operation for a $B$ which falls within a slice that was not allocated before. We need to specify the behavior in this case.
@@
    It is a $\mathtt{SflcWrite}$ operation for a $B$ which falls within a slice that was not allocated before. We need to define how to allocate a new physical slice.
Notice that the offset of a block within a slice is left unchanged by the position map. So, we can easily find the LSI $\sigma$ to which $B$ belongs as simply $\sigma:=\lfloor B/S_{L}\rfloor$. Then we need to check whether $\mathtt{GetSliceMap}(i,\sigma)$ is defined or not. Luckily, as we have seen above, position maps are not too large. So, when decrypting the position maps during device instantiation, we can store in memory a full view of these position maps for each volume as arrays of PSIs indexed by the related LSI, and define a special return symbol $\bot$ for those LSIs which have not yet been assigned.
Then let’s analyse the three cases above one by one. In the first case, we just need to consider the PSI $\varPsi$ just obtained and add a correct offset. This will give us the physical address $B_{\mathsf{phys}}$ where we will find the data to decrypt. From what we have just discussed, this can be done as $B_{\mathsf{phys}}:=\varPsi\cdot S_{P}+\Delta_{S}+(B\mod S_{L})$.
@@
In the third and last case we need a way to allocate a new slice for volume $V_{i}$ at a position consistent with $B$. We do this through a function $\mathtt{CreateSliceMap}$ which returns a PSI uniformly at random among those not yet mapped. We will see in the next paragraph how to implement this function.
Algorithm 1 SflcRead($V_{i},B$)
1:$\sigma:=\lfloor B/S_{L}\rfloor$
2:$\varPsi\leftarrow\mathtt{GetPosMap}(V_{i},\;\sigma)$
3:if $\varPsi=\perp$
@@
5:$B_{\mathsf{phys}}:=\varPsi\cdot S_{P}\,+\,\Delta_{S}\,+\,(B\mod S_{L})$
6:$\mathsf{IV}\leftarrow\mathtt{LoadIV}(B_{\mathsf{phys}})$
7:$d\leftarrow\mathtt{bRead}(B_{\mathsf{phys}})$
8:return $\mathtt{Decrypt}(d,V_{i}.K;\mathsf{IV})$
Algorithm 2 SflcWrite($V_{i},B,d$)
1:$\sigma:=\lfloor B/S_{L}\rfloor$
2:$\varPsi\leftarrow\mathtt{GetPosMap}(V_{i},\;\sigma)$
3:if $\varPsi=\perp$
4:     $\varPsi\leftarrow\mathtt{NewSlice}(V_{i},\;\sigma)$
5:     if $\varPsi=\perp$ return Error
6:$B_{\mathsf{phys}}:=\varPsi\cdot S_{P}\,+\,\Delta_{S}\,+\,(B\mod S_{L})$
7:$\mathsf{IV}\leftarrow\mathtt{SampleAndStoreIV}(B_{\mathsf{phys}})$
[…]
9:return $\mathtt{bWrite}(B_{\mathsf{phys}},\;c)$
@@
Slice mappings are created lazily, only when the first request for a block belonging to a new, yet-unmapped slice arrives. At this point, we create the mapping for this slice by sampling a physical slice uniformly at random among the free ones: this guarantees that no conflicts arise between volumes, and their slices end up randomly interleaved on the disk. To make this lazy sampling possible, we need to implement efficiently a function NewSlice: given as input a volume’s identifier $V_{i}$ and an LSI $\sigma$, it returns a corresponding PSI $\varPsi$ for the new slice. There are different ways to implement this, here we give a reference description using a permutation of the array representing the slices.
@@
Concretely, we keep an in-memory, per-device array $\mathtt{prmslices}$ of PSIs of size $\lfloor\frac{N}{S_{P}}\rfloor$ (the maximum possible number of PSIs), and a bit-array of the same size bfld, an “occupation bitfield” telling us which physical slices are occupied (the PSIs are the indexes of the array). Initially, at device instantiation, $\mathtt{prmslices}$ is initialized as $\mathtt{prmslices}[i]:=i$ and then permuted using the efficient Fisher-Yates algorithm [13]. The bitfield bfld is initialized with all free elements. When a volume is opened, and we discover the physical slices it maps to, we mark them as occupied in bfld. The slice allocation algorithm then simply works by repeatedly taking the next element from the pre-shuffled array of PSIs until the bitfield tells us it is free (at which point we take it and mark it as occupied). A stateful occupation counter $\mathtt{octr}$ is kept to facilitate this task, initially set to $1$, and then increased up to the first element of $\mathtt{prmslices}$ marked as $\mathtt{free}$ (i.e., the first $\mathtt{octr}$ elements of $\mathtt{prmslices}$ are guaranteed to be $\mathtt{occupied}$ in bfld). This way, when a new slice is required, it is immediate to advance to the next available one. We also need to update $V_{i}$’s position map $\mathtt{PosMap}_{i}$ before returning (this is the only change that will eventually persist, encrypted, on disk).

### v2

@@ 2.3.3 Single-snapshot security
A PD scheme $\Pi$ is single-snapshot (SS)-secure iff for any PPT adversary $\mathcal{A}$ which chooses $N\in\left\{2,\ldots,N_{\mathsf{MAX}}\right\}$, passwords $P_{0},\ldots,P_{N-2}$, and access patterns $O_{0}$ and $O_{1}$ subject to the constraints outlined in Section 2.3.2, it holds:
@@
  • $D_{0}$ is the disk snapshot resulting from the application of $O_{0}$ to $\Sigma_{0}$, where $\Sigma_{0}\leftarrow\Pi.\mathtt{setup}(P_{0},\ldots,P_{N-1})$ and $P_{N-1}\raisebox{-1.0pt}{$\,\xleftarrow{\raisebox{-1.0pt}{$\scriptscriptstyle\$$}}\,$}\{0,1\}^{\lambda}$.
  • $D_{1}$ is the disk snapshot resulting from the application of $O_{1}$ to $\Sigma_{1}$, where $\Sigma_{1}\leftarrow\Pi.\mathtt{setup}(P_{0},\ldots,P_{N-2})$.
@@ 4 The Shufflecake Scheme
In this section we present our Shufflecake scheme, we explain its way of operation, and we provide a security analysis. We do not provide concrete parameters in this section, as we want to present a generic framework that can be adapted to various implementations, but we provide our own choice of parameters in Section 5. In particular, in this section we will treat security as asymptotic, without targeting a concrete bit-security level. We assume a target storage block device of $S_{\mathsf{D}}$ blocks, each of them of size $S_{\mathsf{B}}$ bits, and a maximum of $N_{\mathsf{MAX}}$ supported volumes per device.
@@ 4.1 Design
    Initialize a device: this is done only once, when a new device is first prepared for use in Shufflecake. It consists in overwriting the device with random data, asking the user to provide the number $N$ of desired volumes and related passwords, and creating an encrypted header with metadata using this information.
@@
    Instantiate a device: this is the preliminary stage of preparing a Shufflecake-initialized device for use. It consists of reading the user’s provided password, trying to decrypt the device’s header metadata with the derived key, and, if successful, recover information on the available volumes provided.
@@
    Open a volume: using the correctly derived volume key, volume-specific metadata is read from the relevant header section. This metadata is used to create a logical device which is presented to the user, and the user’s OS can issue $\mathtt{SflcRead}$ and $\mathtt{SflcWrite}$ requests to this logical volume.
@@
    Close a previously instantiated device: ephemeral state changes, if present, are written (encrypted) to disk, and then all the open volumes provided by that device are removed from the user’s view.
At its core, Shufflecake is a block indirection layer on top of an encryption layer. Our indirection layer realizes a mechanism which is already a strong improvement over TrueCrypt, since it fixes two of its crucial limitations: it allows multiple volumes, and it is filesystem-independent. Data decryption keys for every volume are derived by a password and other header-derived randomness. Furthermore, the decrypted payload of the header of volume $V_{i}$ with $i>0$ also contains a copy of the header decryption key for volume $V_{i-1}$. This allows to recursively open all volumes present in a device by using a single password, which in turn improves security and user experience, as we will see.
@@
The device’s physical storage space is statically divided into a header section and a data section. The header section is found at the beginning of the disk, and it is composed of a fixed-size device master block (DMB), and $N_{\mathsf{MAX}}$ equal-sized volume headers (each of them comprised of a volume master block (VMB) and a slice map), irrespective of how many volumes there are effectively. This mild waste of space is necessary in order to to prevent the adversary from trivially deducing the number of volumes by the size of the device header (which might be possible by analysing the data allocation pattern even when not all volumes are opened). Let us analyse all these sections, starting with the data.
@@
We distinguish between logical slices, of size $S_{\mathsf{L}}$ blocks, and physical slices, of size $S_{\mathsf{P}}=S_{\mathsf{L}}+\Delta_{\mathsf{S}}$, where $\Delta_{\mathsf{S}}$ blocks are used to store the encryption IVs for that slice. Logical slices store the data of their respective volumes, while physical slices are used to reserve space on the disk to allocate the encrypted logical slice. A physical slice can either be unallocated (i.e., unclaimed by any volume), or mapped to a single logical slice belonging to some volume $V_{i}$. In the latter case, the mapping from the logical slice index (LSI) $\sigma$ of volume $V_{i}$ to the device’s physical slice index (PSI) $\psi$ is given by a correspondence $\mathsf{SliceMap}_{i}:\sigma\mapsto\psi$ which is just a lookup of $V_{i}$’s slice map (basically, a per-volume array). I/O operations to a logical block address $\beta_{\mathsf{L}}$ of volume $V_{i}$ are performed through the two interfaces $\mathtt{SflcRead}$ and $\mathtt{SflcWrite}$. We will describe later these two interfaces, as well as the structure of the slice map.
@@
The DMB encapsulates all password-related data. It begins with one single KDF salt, shared for all volumes: this salt, combined with a volume $V_{i}$’s password through a KDF, yields the volume’s key-encryption-key ($\mathsf{KEK}_{i}$). Notice that we derive every $\mathsf{KEK}_{i}$ by using just a single global salt for all $N_{\mathsf{MAX}}$ volumes, otherwise we would incur in up to $N_{\mathsf{MAX}}$ expensive different key derivations every time we instantiate a device [^5] This limitation can actually be avoided by using some key derivation tricks, like ‘re-salting’, i.e. using the output of the KDF in combination with a (fast) hash using a per-header salt. We leave this approach for future consideration.. This does not hamper security, as password hashes are never stored on disk, but only used to generate the KEK which is in turn used as a decryption key.
Then come $N_{\mathsf{MAX}}$ DMB cells, each being an authenticated ciphertext (together with the corresponding IV), encrypted with the respective volume’s KEK: the plaintext is itself another cryptographic key, encrypting the volume’s VMB (the volume master key $\mathsf{VMK}_{i}$). This key decoupling allows us to possibly change a volume’s password without having to re-encrypt all its content with a different key. For granularity and consistency, the overall size of the DMB is fixed to be exactly one block. The rest of the DMB, and the unused DMB cells, contain random noise.
@@
The header of volume $V_{i}$ is composed of a Volume Master Block (VMB), followed by the encrypted (but not authenticated) slice map of $V_{i}$. We discuss the slice map in the next paragraph. The VMB is a single block containing a (non-authenticated) ciphertext, encrypted with the volume master block’s key $\mathsf{VMK}_{i}$, and the associated IV. The underlying plaintext is composed of the following fields:
  • The volume’s volume-encryption key ($\mathsf{VEK}_{i}$), used for encrypting the actual data section and the slice map.
  • The previous volume’s VMB key $\mathsf{VMK}_{i-1}$ (or a random value if $i=0$).
  • The number of slices $\mathsf{numslices}$ contained in the device.
@@
The device-specific value $\mathsf{numslices}$ defines and fixes the size of the slice maps, even in the case that the device is resized; it is replicated across all volume headers in order to be decryptable with any provided password. The presence of $\mathsf{VMK}_{i-1}$ is what allows us to impose a hierarchy on the otherwise-independent volumes (they are all treated equally in the data section). This way, once we open volume $V_{i}$, we can iteratively walk the backwards-linked list induced by this field, and also open volumes $V_{i-1}$ through $V_{0}$. While this approach compromises deniability for a volume $V_{i}$ “in the middle”, once the password to $V_{j}\,,\,j>i$ has been provided, it does not harm deniability as defined by the security game: the volume we want to hide is the last one, $V_{N-1}$, not the middle ones. The usefulness of this approach is discussed in Section 4.2.
@@
Slice slice maps are arrays $\mathsf{SliceMap}_{i}$ of $\mathsf{numslices}$ elements, where every element is a PSI: the index of each element is the LSI mapping to that PSI. Each volume’s slice map is decrypted and loaded in memory when the volume is instantiated; these maps are kept entirely in RAM while the volumes are “live”, and persisted on-disk (encrypted, together with their fresh IVs) when the volumes are closed. slice maps are stored after the VMBs as equal-size ciphertexts, large enough to address all the possible physical slices. The RAM and disk space requirements are modest: if the underlying device is $S_{\mathsf{D}}$ blocks large and the header size is $S_{\mathsf{H}}$ blocks, there can be at most $\mathsf{mxslc}=\lceil\frac{S_{\mathsf{D}}-S_{\mathsf{H}}}{S_{\mathsf{P}}}\rceil$ physical slices, each requiring $O(\log\frac{S_{\mathsf{D}}}{S_{\mathsf{P}}})$ bits to be indexed, for a total size of $O(\frac{S_{\mathsf{D}}}{S_{\mathsf{P}}}\log\frac{S_{\mathsf{D}}}{S_{\mathsf{P}}})$ bits per slice map. This is in turn due to the choice of addressing the storage space at the slice granularity, instead of the block granularity, which would have entailed a slice map of $O(S_{\mathsf{D}}\log S_{\mathsf{D}})$ bits.
@@
We now describe more in detail how Shufflecake operates and explain the rationale behind certain design choices. We start by explaining how we use encryption to secure data on-disk. Then we look at the indirection layer between physical data on disk and data on the logical volumes, i.e., the mechanism which creates the correspondence between physical and logical slices. We explain how this correspondence is updated when new logical slices are needed as data is written on the logical volumes. Then, we explain how Shufflecake handles the case of allocation errors, i.e., when a physical slice is mistakenly assigned to more than one volume. Finally, we detail how the instantiation procedure works.
@@
In our case, however, we stick to explicitly random IVs because we want to keep the future possibility to extend Shufflecake with some degree of multi-snapshot security, requiring us to re-encrypt some blocks with a different IV while leaving their content unchanged. For this reason, we use CTR mode instead, and the IV of a block is refreshed at each $\mathtt{SflcWrite}$ for that block, to avoid IV-reuse attacks. This means that all these IVs are stored on-disk.
This strategy introduces a potential performance issue: a naive implementation would translate each logical $\mathtt{SflcWrite}$ to a physical $\mathtt{bWrite}$ of the corresponding physical data block, plus an additional bRead-update-bWrite of the whole corresponding IV block. This would be very wasteful in terms of I/O overhead, because we only need to update one IV (e.g., 16 bytes for AES-CTR), but we are forced to load and store whole blocks (typically 4096 bytes).
We avoid this problem by caching IV blocks in RAM, in an LRU cache of predefined depth (e.g., 1024 entries). For the performance reasons just discussed, this cache is not write-through. This way, we coalesce possibly many updates of the same IV block (triggered by many logical $\mathtt{SflcWrite}$s to the same data block, or by logical $\mathtt{SflcWrite}$s to many blocks within the same slice) into just one physical $\mathtt{bWrite}$, thereby lowering the I/O overhead.
For each physical slice, we pack the IVs for the blocks contained therein into $\Delta_{\mathsf{S}}$ physical blocks at the beginning of the slice. There is a simple static correspondence between a physical block and the on-disk location of its IV: the $(j+\Delta_{\mathsf{S}})$-th block within a physical slice is encrypted by the $j$-th IV within the initial $\Delta_{\mathsf{S}}$ blocks. Hence, we assume the existence of functions $\mathtt{LoadIV}$ and $\mathtt{SampleAndStoreIV}$ (with self-explanatory behaviour) which take as input a physical address $\beta_{\mathsf{P}}$ and return the corresponding IV. Analogously, we consider the functions $\mathtt{Encrypt}({p},\mathsf{IV},{k})$ and $\mathtt{Decrypt}({c},\mathsf{IV},{k})$ as acting on blocks according to the implied mode of operation.
@@
Consider a read or write operation to a logical block address $\beta_{\mathsf{L}}$ for volume $V_{i}$. There are three possible cases:
@@
    The requested operation (read or write) happens on a block whose logical address $\beta_{\mathsf{L}}$ was previously allocated. In this case we need to map efficiently $\beta_{\mathsf{L}}$ for volume $V_{i}$ to the corresponding physical address $\beta_{\mathsf{P}}$ on the device.
@@
    It is a $\mathtt{SflcRead}$ operation for a $\beta_{\mathsf{L}}$ which falls within a slice that was not allocated before. We need to specify the behavior in this case.
@@
    It is a $\mathtt{SflcWrite}$ operation for a $\beta_{\mathsf{L}}$ which falls within a slice that was not allocated before. We need to define how to allocate a new physical slice.
Notice that the offset of a block within a slice is left unchanged by the slice map. So, we can find the LSI $\sigma$ to which $\beta_{\mathsf{L}}$ belongs as simply $\sigma:=\lfloor\beta_{\mathsf{L}}/S_{\mathsf{L}}\rfloor$. Then we need to check whether $\mathsf{SliceMap}_{i}[\sigma]$ is defined or not. Luckily, as we have seen above, slice maps are not too large. So, when decrypting the slice maps during device instantiation, we can store in memory a full view of these slice maps for each volume as arrays of PSIs indexed by the related LSI, and define a special return symbol $\bot$ for those LSIs which have not yet been assigned.
Then let’s analyse the three cases above one by one. In the first case, we just need to consider the PSI $\psi$ just obtained and add a correct offset. This will give us the physical address $\beta_{\mathsf{P}}$ where we will find the data to decrypt. From what we just discussed, this can be done as $\beta_{\mathsf{P}}:=\psi\cdot S_{\mathsf{P}}+\Delta_{\mathsf{S}}+(\beta_{\mathsf{L}}\mod S_{\mathsf{L}})$.
@@
In the third and last case we need a way to allocate a new slice for volume $V_{i}$ at a position consistent with $\beta_{\mathsf{L}}$. We do this through a function $\mathtt{NewSlice}$ which returns a PSI uniformly at random among those not yet mapped. We will see in the next paragraph how to implement this function.
Algorithm 1 $\mathtt{SflcRead}$ ($V_{i},\beta_{\mathsf{L}}$)
1:$\sigma:=\lfloor\beta_{\mathsf{L}}/S_{\mathsf{L}}\rfloor$
2:$\psi\leftarrow\mathsf{SliceMap}_{i}[\sigma]$
3:if $\psi=\bot$
@@
5:$\beta_{\mathsf{P}}:=\psi\cdot S_{\mathsf{P}}+\Delta_{\mathsf{S}}+(\beta_{\mathsf{L}}\mod S_{\mathsf{L}})$
6:$\mathsf{IV}\leftarrow\mathtt{LoadIV}(\beta_{\mathsf{P}})$
7:${c}\leftarrow\mathtt{bRead}(\beta_{\mathsf{P}})$
8:return $\mathtt{Decrypt}({c},\mathsf{IV},\mathsf{VEK}_{i})$
Algorithm 2 $\mathtt{SflcWrite}$ ($V_{i},B,d$)
1:$\sigma:=\lfloor\beta_{\mathsf{L}}/S_{\mathsf{L}}\rfloor$
2:$\psi\leftarrow\mathsf{SliceMap}_{i}[\sigma]$
3:if $\psi=\bot$
4:     $\psi\leftarrow\mathtt{NewSlice}(V_{i},\sigma)$
5:     if $\psi=\bot$ return Error
6:$\beta_{\mathsf{P}}:=\psi\cdot S_{\mathsf{P}}+\Delta_{\mathsf{S}}+(\beta_{\mathsf{L}}\mod S_{\mathsf{L}})$
7:$\mathsf{IV}\leftarrow\mathtt{SampleAndStoreIV}(\beta_{\mathsf{P}})$
8:${c}:=\mathtt{Encrypt}(d,\mathsf{IV},\mathsf{VEK}_{i})$
9:return $\mathtt{bWrite}(\beta_{\mathsf{P}},{c})$
@@
slice mappings are created lazily, only when the first request for a block belonging to a new, yet-unmapped slice arrives. At this point, we create the mapping for this slice by sampling a physical slice uniformly at random among the free ones: this guarantees that no conflicts arise between volumes, and their slices end up randomly interleaved on the disk. To make this lazy sampling possible, we need to implement efficiently a function $\mathtt{NewSlice}$: given as input a volume’s identifier $V_{i}$ and an LSI $\sigma$, it returns a corresponding PSI $\psi$ for the new slice. There are different ways to implement this, here we give a reference description using a permutation of the array representing the slices.
@@
Concretely, we keep an in-memory, per-device array $\mathsf{prmslices}$ of PSIs of size at most $\mathsf{mxslc}=\lceil\frac{S_{\mathsf{D}}-S_{\mathsf{H}}}{S_{\mathsf{P}}}\rceil$ (maximum possible number of slices), and an array of the same size $\mathsf{ofld}$, an “occupation bitfield” telling us which physical slices are occupied (the PSIs are the indexes of the array). Initially, at device instantiation, $\mathsf{prmslices}$ is initialized as $\mathsf{prmslices}[i]:=i$ and then permuted using the efficient Fisher-Yates algorithm [13]. The bitfield $\mathsf{ofld}$ is initialized with all $\mathsf{free}$ values. When a volume is $\mathtt{open}$ed, and we discover the physical slices it maps to, we mark them as $\mathsf{occupied}$ in $\mathsf{ofld}$. The slice allocation algorithm then simply works by repeatedly taking the next element from the pre-shuffled array of PSIs until the bitfield tells us it is free (at which point we take it and mark it as occupied). A stateful occupation counter $\mathsf{octr}$ is kept to facilitate this task, initially set to 00, and then increased up to the first element of $\mathsf{prmslices}$ marked as $\mathsf{free}$ (i.e., the first $\mathsf{octr}$ elements of $\mathsf{prmslices}$ are guaranteed to be $\mathsf{occupied}$ in $\mathsf{ofld}$). This way, when a new slice is required, it is immediate to advance to the next available one. We also need to update $V_{i}$’s slice map $\mathsf{SliceMap}_{i}$ before returning (this is the only change that will eventually persist, encrypted, on disk).
Algorithm 3 $\mathtt{NewSlice}$ ($V_{i},\sigma$)
1:while $\mathsf{ofld}[\mathsf{prmslices}[\mathsf{octr}]]=\mathsf{occupied}$ do: $\mathsf{octr}:=\mathsf{octr}+1$
2:$\psi_{\mathsf{new}}\leftarrow\mathsf{prmslices}[\mathsf{octr}]$
3:$\mathsf{ofld}[\mathsf{prmslices}[\mathsf{octr}]]:=\mathsf{occupied}$
4:$\mathsf{SliceMap}_{i}[\sigma]:=\psi_{\mathsf{new}}$
5:return $\psi_{\mathsf{new}}$